a ball is thrown upward into the air with an intitial velocity of 64 feet per second.the formula h(t)= 64t-16t^2 gives it height above the ground after t seconds.
a) what is the height after 1.5 seconds?
b) what is the maximum height?
c) how many seconds will pass before it returns to the ground?

Answer by lwsshak3(11628)

a ball is thrown upward into the air with an intitial velocity of 64 feet per second.the formula h(t)= 64t-16t^2 gives it height above the ground after t seconds.
..
a) what is the height after 1.5 seconds?
h(t)= 64t-16t^2
h(1.5)=64*1.5-16*1.5^2=60
ans: height after 1.5 seconds=60 feet
b) what is the maximum height?
complete the square
h(t)=-16(t^2-4t+4)+64
h(t)=-16(t-2)^2+64
ans:
maximum height=64 feet
c) how many seconds will pass before it returns to the ground?
ans:4 sec